how do I solve 2c + 3 = 4 (c -1) and 9 - (3r - 1) = 12? please help quick
Answer by josgarithmetic(39623)

Solve those?
I will do the second equation but not to completion.
9-(3r-1)=12
-3r+1=12-9
3r-1=-(12-9)
3r=1+(-12)+9
Can you finish?